• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name paschke

    Log entry time 22:44:47 on February 4,2003

    Entry number 94840

    keyword=HAPPEX DAQ activity



    HAPPEX DAQ activity for the evening included
    - IA scan
    - Sign Check, (N2 target)
    - Carbon target, with and w/o OS
    - DAC noise OFF test (w/ & w/o OS)
    - Lumis checked in another ADC
    - bcm3 (bcm2 - 3X output) moved to ADC0_1

    In the following OS1 means:
    OS=0, IT=13300, All counting house ADCs are in Low gain except ADC0 & 4, which are High Gain
    while OS13 means
    OS=12, IT=1000, All counting house ADCs are in High gain


    CH Inj Comments
    1703 long "pedestal" run, from 11:30-15:00.
    5227 "pedestal" run (no forced gain)
    1704 "ped" run with OS=12, IT=1000, All ADC hi-gain (300 k events)
    1705 "ped" run with OS=12, IT=1000, All ADC hi-gain (877 k events, whoops)
    1706 "ped" run with OS=14, IT=100, ADC=HI... beam came...

    1707 5uA, OS=13 (14) (whoops)
    1708 5228 4uA, IT=1000, OS=12, ADC=HiGain
    **********IA Scan********************
    1709 5229 IA=4.14, OS=1 (IT=13300, ADC=LoGain, except 0&4=HI) HWP=OUT, RWP=6435
    1710 5230 IA=9.00, OS=1 (IT=13300, ADC=LoGain, except 0&4=HI)
    1711 5231 IA=1.00, OS=1 (IT=13300, ADC=LoGain, except 0&4=HI)
    1712 5232 IA=5.00, OS=1 (IT=13300, ADC=LoGain, except 0&4=HI)
    **************************************
    **********Sign Check******************
    1713 5233 N2 reference cell filling, started at 100 psi, 5uA
    "POS" Pita voltage turned OFF (setpoint = 0.000, readback = 0.04) (Neg = 7.752, readback 2.07)
    1714 5234 N2 cell, 149psig, 12 uA (lumi2 over range)
    1715 5235 N2 cell, 149psig, 11 uA
    1716 5236 OS13 config, N2 cell, 141psig, 11 uA
    IHWP inserted, RWP left at 6435
    1717 5237 OS1 config, N2 cell 147 psig
    1718 5238 OS13 config
    1719 5239 OS1 config
    "POS" Pita turned back on, (setpt=7.690, readback = 3.25)
    **************************************
    Carbon target, raster off
    1720 5240 OS1
    1721 | OS13
    1722 | OS13, Lumi HV = Nominal + 100V
    1723 | OS13, Lumi HV = Nominal + 200V
    Lumi HV back to nominal, Pol 3He target
    RHWP set to 6650 (nominal for IHWP=IN)
    1724 OS1, feedback ok.
    Feedback disabled (18:33) to take oversampled data
    1725 OS13 Lumi HV= nominal
    1726 OS13 Lumi HV= nominal + 100 V
    1727 OS13 Lumi HV= nominal + 200 V
    1728 OS1, Feedback on and OK. Killed badly (no end-of-run record)
    1729 OS1, Feedback on and OK.
    **********DAC Noise Study*************
    1730 DAC Noise Turned off, OS1 (Feedback on and OK)
    1731 DAC Noise Turned off, OS13 (feedback disabled)
    1732 DAC Noise Turned off, OS1 (Feedback on and OK)
    **************************************
    **********Moved Lumi to new ADC*******
    Moved Lumi channels to corresponding channels (0, 1) on ADC 1
    Moved timing ribbon cable to this ADC also. Modified template for this.
    1733 DAC Noise OFF, ended with a beam trip (good for pedestals?)
    1734 DAC Noise OFF, same thing.
    1735 DAC Noise OFF, new template to correct bcm3 position (ADC4_0, not ADC4_1)
    1736 DAC noise OFF, moved bcm3 to ADC0_1
    Moved Lumi channels (lumi1 and lumi2) back to ADC2_0 and ADC2_1 respectively.
    Moved timing ribbon cable back to this ADC also. Modified template for this.
    1737 DAC Noise back on. Whoops, control.db not changed.
    1738 DAC Noise ON. Control.db is correct. Feedback is on and OK.
    **************************************